# Dynamic Prediction for Hospital Readmission in Patients with Chronic Heart Failure

### v1

@@ Results
In parallel, Figure 3 summarizes the estimated associations between baseline covariates and the composite outcome using the Weibull survival model. The forest plot illustrates hazard ratios with 95% confidence intervals, providing a benchmark static risk profile for the cohort. Each hazard ratio represents the multiplicative change in the risk of the event associated with a one-unit increase (or presence of a condition), holding all other predictors constant. Several predictors emerge as statistically significant drivers of adverse outcomes, including BMI, creatinine, systolic blood pressure, atrial fibrillation, anemia, chronic kidney disease, COPD, and hypertension. These associations align with established clinical understanding of HF prognosis [3] and confirm that the baseline model captures clinically meaningful patterns before incorporating longitudinal NT-proBNP information.
@@
Individualized dynamic predictions
@@
These findings demonstrate that the full trajectory of NT-proBNP provides critical prognostic information that is lost when using only static, single-timepoint methods. This is a key contribution of our work and stands in contrast to previous studies [26], which concluded that the last observed NT-proBNP value may suffice for risk prediction. Our results challenge this, highlighting that risk is not static but evolves. As shown in our performance analysis, the joint model’s ability to dynamically incorporate the direction of the biomarker trend (i.e., improving or worsening) provides a more accurate and adaptive risk assessment than a static benchmark model. Our model therefore supports a more sophisticated clinical decision-making process, informed by the patient’s entire biomarker trajectory rather than a single, isolated snapshot.
Predictive performance
@@ Discussion
A key strength of our work is the demonstration that dynamic prediction provides clinically relevant insights that static models cannot recover. For many patients, NT-proBNP trends during the early post-discharge period reflected meaningful changes in risk that were invisible when using a single measurement. Patients with decreasing trajectories showed correspondingly lower predicted risk, whereas increasing trajectories signaled elevated risk. This ability to update outcome predictions based on dynamic changes of the biomarker is one of the major advantages of joint modeling and highlights its potential utility in routine heart failure management. Our findings supplement and clarify prior work in the field: while prior research [26] suggested that the most recent NT-proBNP value might be sufficient for prediction, our results indicate that incorporating full trajectories leads to more accurate and adaptive risk estimates, particularly when additional follow-up information becomes available.
@@
Data Availability
@@ Acknowledgements
Author contributions statement
S. Metelli and F. Mercier conceptualized and designed the project. R. Farina wrote the code, performed all analyses, and drafted the manuscript. C. Wohlfart contributed to data curation, analyses, and interpretation of data. S. Masson contributed to clinical interpretation and relevance. All authors revised the manuscript and accepted the final version.
Competing interests
S. Masson is an employee at Roche Diagnostics, that commercializes an NT-proBNP assay. All other authors declare no conflicts of interest.

### v2

@@ 3.2 Model fit and baseline risk associations
In parallel, Figure 3 summarizes the estimated associations between baseline covariates and the composite outcome using the Weibull survival model. The forest plot illustrates hazard ratios with 95% confidence intervals, providing a benchmark static risk profile for the cohort. Each hazard ratio represents the multiplicative change in the risk of the event associated with a one-unit increase (or presence of a condition), holding all other predictors constant. Several predictors emerge as statistically significant drivers of adverse outcomes, including BMI, creatinine, systolic blood pressure, atrial fibrillation, anemia, chronic kidney disease, COPD, and hypertension. These associations align with established clinical understanding of HF prognosis [ponikowski20162016] and confirm that the baseline model captures clinically meaningful patterns before incorporating longitudinal NT-proBNP information.
@@
3.3 Individualized dynamic predictions
@@
These findings demonstrate that the full trajectory of NT-proBNP provides critical prognostic information that is lost when using only static, single-timepoint methods. This is a key contribution of our work and stands in contrast to previous studies [zhang2018], which concluded that the last observed NT-proBNP value may suffice for risk prediction. Our results challenge this, highlighting that risk is not static but evolves. As shown in our performance analysis, the joint model’s ability to dynamically incorporate the direction of the biomarker trend (i.e., improving or worsening) provides a more accurate and adaptive risk assessment than a static benchmark model. Our model therefore supports a more sophisticated clinical decision-making process, informed by the patient’s entire biomarker trajectory rather than a single, isolated snapshot.
3.4 Predictive performance
@@ 4 Discussion
A key strength of our work is the demonstration that dynamic prediction provides clinically relevant insights that static models cannot recover. For many patients, NT-proBNP trends during the early post-discharge period reflected meaningful changes in risk that were invisible when using a single measurement. Patients with decreasing trajectories showed correspondingly lower predicted risk, whereas increasing trajectories signaled elevated risk. This ability to update outcome predictions based on dynamic changes of the biomarker is one of the major advantages of joint modeling and highlights its potential utility in routine heart failure management. Our findings supplement and clarify prior work in the field: while prior research [zhang2018] suggested that the most recent NT-proBNP value might be sufficient for prediction, our results indicate that incorporating full trajectories leads to more accurate and adaptive risk estimates, particularly when additional follow-up information becomes available.
@@ 5 Conclusion
Supplementary Information
Abbreviations
AFib Atrial Fibrillation
BMI Body Mass Index
CKD Chronic Kidney Disease
COPD Chronic Obstructive Pulmonary Disease
HF Heart Failure
IBS Integrated Brier Score
ICI Integrated Calibration Index
JM Joint Model / Joint Modeling
LVEF Left Ventricular Ejection Fraction
MCMC Markov Chain Monte Carlo
NT-proBNP N-terminal pro-B-type natriuretic peptide
NYHA New York Heart Association
SBP Systolic Blood Pressure
@@ Acknowledgements
Availability of data and materials